OpenStudy (anonymous):

Marco makes $36,000 a year. If he receives a 5% raise, how much more money will he be earning each year? $ ________

OpenStudy (aaronandyson):

First:Calculate 5% of 36,000 Second:Add the above answer to 36,000

OpenStudy (aaronandyson):

Solution for what is 5% of 36000 36000/x=100/5 (36000/x)*x=(100/5)*x - we multiply both sides of the equation by x 36000=20*x - we divide both sides of the equation by (20) to get x 36000/20=x

OpenStudy (aaronandyson):

x = ?